Is x = 3 an asymptote of the function f(x)= (x[sup]2[/sup] -5x+6)/(x-3)? Why?
1
Expert's answer
2011-05-19T08:16:41-0400
f(x) = (x2 - 5x + 6)/(x-3)
Let's factorize the numerator:
x2 - 5x + 6 = (x-3)(x-2)
Thus f(x) = (x-3)(x-2)/(x-3) = x-2.
Therefore x=3 is not the asymptote.